### v1

# Distributed games with jumps: an $\alpha$-potential game Approach

Abstract.
Motivated by game-theoretic models of crowd motion dynamics, this paper analyzes a broad class of distributed games with jump diffusions within the recently developed $\alpha$-potential games framework. We demonstrate that analyzing the $\alpha$-Nash equilibria is equivalent to solving finite-dimensional control problems. Beyond the viscosity and verification characterizations for the general games, we explicitly and in detail examine how spatial population distributions and interaction rules influence the structure of $\alpha$-Nash equilibria in these distributed settings, and in particular for crowd motion games.
Our theoretical results are supported by numerical implementations using policy gradient-based algorithms, further demonstrating the computational advantages of the $\alpha$-potential game framework in computing Nash equilibria for general dynamic games.
Key words and phrases:
Potential game, Nash equilibrium, distributed game, Hamilton–Jacobi–Bellman equation, policy-gradient algorithm, game-theoretic motion planning
2020 Mathematics Subject Classification:
91A06, 91A15, 91A14, 68T07
1. Introduction
Motivating example and distributed games.
Consider the following motion planning game [27, 3, 39, 7], where a group of $N$ players each controls or chooses their preferred route to reach their respective destinations; their paths are impacted by the spatial distribution of the population and their interactions. In this game, each player aims to find the optimal path according to her cost functional consisting of terminal costs and the running costs which depend on the controls and the path to her destination. This game can be modeled as the following stochastic differential game. For each player $i\in[N]\coloneqq\{1,2,\cdots,N\}$, given her control process $u_{i}$, her state process $X^{u_{i}}_{i}$, representing the player’s position, is governed by the following controlled jump-diffusion process:
${\mathrm{d}}X_{i,t}$ $=b_{i}(t)u_{i,t}{\mathrm{d}}t+\sigma_{i}(t){\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}\gamma_{ij}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad t\in(0,T];\quad X_{i,0}=x_{i}\in{\mathbb{R}}^{d},$ (1.1)
where $b_{i}:[0,T]{{\rightarrow}}{\mathbb{R}}^{d\times k}$, $\sigma_{i}:[0,T]{{\rightarrow}}{\mathbb{R}}^{d\times n}$ and $\gamma_{ij}:[0,T]\times{\mathbb{R}}^{p}{{\rightarrow}}{\mathbb{R}}^{d}$ are measurable functions such that (1.1) admits a unique strong solution on an appropriate probability space which supports the $n$-dimensional Brownian motion $W$ and the jump processes $(\tilde{\eta}_{j})_{j=1}^{m}$. Given a joint control profile ${\textbf{{u}}}=(u_{i})_{i\in[N]}$ from an admissible set, each player $i$ aims to minimize over her admissible controls an objective function of the form
$J_{i}({\textbf{{u}}})\coloneqq{\mathbb{E}}\left[\int_{0}^{T}\left(\ell_{i}(u_{i,t})+\frac{1}{N-1}\sum_{j=1,j\neq i}^{N}q_{ij}K(X^{u_{i}}_{i,t}-X^{u_{j}}_{j,t})\right)\,{\mathrm{d}}t+g_{i}(X^{u_{i}}_{i,T})\right],$ (1.2)
where $\ell_{i}$ is the cost of control and $g_{i}$ is the terminal cost. The kernel function $K$ can be specified to model self-organizing behavior such as flocking, or aversion behavior, with adjustment of the interaction intensity by $q_{ij}\geq 0$.
The above crowd motion game is a special class of stochastic differential games which we name as distributed games; See Section 2 for the detailed formulation of these games. The term “distributed” refers to the characteristics of the game where each player’s dynamics evolve according to a controlled stochastic process that depends only on her own control, while her objective function may depend on the joint state and control profiles of all players (see also Remark 2.1). Such a framework has been used in a variety of applications where agents interact through their objectives but evolve independently in state, including distributed control of multi-agent systems and trajectory planning [23, 42, 43, 1], transportation and routing [5, 9, 25], as well as in energy markets and smart grids [29, 41, 36, 34, 45, 32].
In general, deriving Nash equilibria for this type of game is analytically challenging, as the interaction kernel $K$ in (1.2) is typically non-convex, which precludes the use of standard tools such as the stochastic maximum principle. An exception arises in the special case of mean field games, under the assumptions that players are homogeneous and interact weakly through empirical measures, and the number of players $N\to\infty$, see for instance, [27, 3, 39, 7].
Meanwhile, the recently introduced $\alpha$-potential game framework has shown significant promise for analyzing and solving general dynamic games, both from theoretical and algorithmic perspectives [15, 18, 17, 16, 22, 30]. For example, [15] and [17] demonstrate that computing a Nash equilibrium can be reformulated as an optimization problem involving a single $\alpha$-potential function and the analysis of the parameter $\alpha$. Furthermore, optimizing this $\alpha$-potential function has been shown in [17] to be equivalent to solving a conditional McKean–Vlasov control problem. Consequently, an $\alpha$-Nash equilibrium of the stochastic game can be characterized by an infinite-dimensional Hamilton–Jacobi–Bellman (HJB) equation. Unlike the conventional mean field game approach, which relies on weak interactions among players or takes the limit as the number of players $N\to\infty$, the $\alpha$-potential game framework can be directly applied to finite-player settings.
Our approach and our work.
In this paper, we adopt the $\alpha$-potential game framework to analyze distributed games with controlled jump diffusions under an open-loop setting. Within this framework, we show that the task of finding an $\alpha$-Nash equilibrium can be further reduced to solving a finite-dimensional control problem (Theorem 3.1). This reduction enables the application of standard tools, such as the viscosity solution method and the verification theorem, to characterize the $\alpha$-Nash equilibria via a finite-dimensional HJB equation (Theorems 4.1 and 4.2).
Moreover, we fully characterize how the parameter $\alpha$ depends on the underlying game characteristics. In particular, for the class of crowd motion games, we explicitly demonstrate how the resulting $\alpha$-Nash equilibria are shaped by the spatial distribution of the population, as well as the intensity and asymmetry of players’ interactions. This dependence is exemplified through the choice of the kernel function $K$ and the structure of the interaction weights $(q_{ij})_{i,j\in[N]}$ in the agents’ payoffs (1.2) (Theorem 6.1 and Corollary 6.1).
Finally, leveraging the algorithmic advantages of the $\alpha$-potential game paradigm as illustrated in [15], we develop an efficient policy-gradient algorithm (Algorithm 1) to minimize the $\alpha$-potential function and thereby construct an $\alpha$-Nash equilibrium. Through numerical experiments on crowd motion games, we showcase distinct emergent trajectories in both flocking and aversion dynamics.
Notation
Let $T>0$. For each measurable function $\phi:[0,T]\to{\mathbb{R}}^{n}$, we define its $L^{2}$-norm $\|\phi\|_{L^{2}}=\left(\int_{0}^{T}|\phi(s)|^{2}{\mathrm{d}}s\right)^{1/2}$ with ||$\cdot$|| ⋅ | being the Euclidean norm, and for each $\phi:U\to{\mathbb{R}}^{m\times n}$ defined on a set $U$, we define its sup-norm $\|\phi\|_{L^{\infty}}=\sup_{u\in U}\|\phi(u)\|_{\textrm{sp}}$, with $\|\cdot\|_{\textrm{sp}}$ being the spectral norm of a matrix.
For each filtered probability space $(\Omega,\mathcal{F},{\mathbb{F}},\mathbb{P})$ and Euclidean space $(E,|\cdot|)$, we denote by $\mathcal{S}^{2}(E)$ the space of $E$-valued ${\mathbb{F}}$-progressively measurable processes $X:\Omega\times[0,T]\to E$ satisfying $\|X\|_{\mathcal{S}^{2}(E)}={\mathbb{E}}[\sup_{s\in[0,T]}|X_{s}|^{2}]^{1/2}<\infty$, and by $\mathcal{H}^{2}(E)$ the space of $E$-valued ${\mathbb{F}}$-progressively measurable processes $X:\Omega\times[0,T]\to E$ satisfying $\|X\|_{\mathcal{H}^{2}(E)}={\mathbb{E}}[\int_{0}^{T}|X_{s}|^{2}{\mathrm{d}}s]^{1/2}<\infty$. With a slight abuse of notation, for any $m,n\in{\mathbb{N}}$, we identify the product spaces $\mathcal{S}^{2}({\mathbb{R}}^{n})^{m}$ and $\mathcal{H}^{2}({\mathbb{R}}^{n})^{m}$ with $\mathcal{S}^{2}({\mathbb{R}}^{mn})$ and $\mathcal{H}^{2}({\mathbb{R}}^{mn})$, respectively.
2. Distributed Games and Their Nash Equilibria
This section introduces a class of stochastic differential games, referred to as distributed games, in which each player’s dynamics evolve according to a drift-controlled jump-diffusion process that depends only on their own control, while their objective function may depend on the joint state and control profiles of all players. We next present preliminary results for applying the $\alpha$-potential game framework developed in [17] to compute approximate Nash equilibria for such games.
2.1. Mathematical Setup
Let $T>0$ be given terminal time, and $N,d,n,m,p\in\mathbb{N}$. Let $(\Omega,\mathcal{F},\mathbb{P})$ be a complete probability space which supports the following three mutually independent processes: a family of square integrable $d$-dimensional random variables $(\xi_{i})_{i=1}^{N}$, an $n$-dimensional Brownian motion $W=(W_{i})_{i=1}^{n}$, and a family of independent Poisson random measures $\eta=(\eta_{i})_{i=1}^{m}$ on $[0,T]\times\mathbb{R}^{p}_{0}$, where $\mathbb{R}^{p}_{0}\coloneqq{\mathbb{R}}^{p}\setminus\{0\}$ is equipped with its Borel $\sigma$-algebra. The random variables $(\xi_{i})_{i=1}^{N}$ represents the initial conditions of the system states, and the processes $W$ and $\tilde{\eta}$ represent the underlying system noises. We assume that each $\eta_{i}$ has a compensator $\nu_{i}({\mathrm{d}}z)\,{\mathrm{d}}t$, with $\nu_{i}$ being a $\sigma$-finite measure on $\mathbb{R}^{p}_{0}$ satisfying $\int_{\mathbb{R}^{p}_{0}}\min(1,|z|^{2})\,\nu_{i}({\mathrm{d}}z)<\infty,$ and define $\widetilde{\eta}_{i}({\mathrm{d}}t,{\mathrm{d}}z)=\eta_{i}({\mathrm{d}}t,{\mathrm{d}}z)-\nu_{i}({\mathrm{d}}z)\,{\mathrm{d}}t$ as the compensated Poisson random measure of $\eta_{i}$. Let $\mathbb{F}=(\mathcal{F}_{t})_{t\in[0,T]}$ be the filtration generated by $(\xi_{i})_{i=1}^{N}$, $W$ and $\eta$, augmented with the $\mathbb{P}$-null sets.
We consider a stochastic differential game involving $N$ players, each employing open-loop control strategies defined as follows. For each $i\in[N]\coloneqq\{1,\ldots,N\}$, let $A_{i}\subset{\mathbb{R}}^{k}$ be player $i$’s action set, and let $\mathcal{A}_{i}$ be the set of player $i$’s admissible controls defined by
$\mathcal{A}_{i}\coloneqq\{u:\Omega\times[0,T]\to A_{i}\mid u\in\mathcal{H}^{2}({\mathbb{R}}^{k})\}.$ (2.1)
Let $A=\prod_{i\in[N]}A_{i}$ be the set of joint action profiles of all players and $\mathcal{A}=\prod_{i\in[N]}\mathcal{A}_{i}$ be the joint control profiles. For each $i\in[N]$, we denote by $\mathcal{A}_{-i}\coloneqq\prod_{j\in[N]\setminus\{i\}}\mathcal{A}_{j}$ the set of control profiles of all players except player $i$, and by ${\textbf{{u}}}=(u_{i})_{i\in[N]}$ and ${\textbf{{u}}}_{-i}=(u_{j})_{j\in[N]\setminus\{i\}}$ a generic element of $\mathcal{A}$ and $\mathcal{A}_{-i}$, respectively.
Given the control sets $(\mathcal{A}_{i})_{i\in[N]}$, each player influences their evolution by controlling the drift of a jump-diffusion process. More precisely, for each ${\textbf{{u}}}=(u_{i})_{i\in[N]}\in\mathcal{A}$, consider the following state dynamics: for all $i\in[N]$,
${\mathrm{d}}X_{i,t}$ $=b_{i}(t)u_{i,t}{\mathrm{d}}t+\sigma_{i}(t){\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}\gamma_{ij}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad t\in(0,T];\quad X_{i,0}=\xi_{i},$ (2.2)
where $b_{i}:[0,T]{{\rightarrow}}{\mathbb{R}}^{d\times k}$, $\sigma_{i}:[0,T]{{\rightarrow}}{\mathbb{R}}^{d\times n}$ and $\gamma_{ij}:[0,T]\times{\mathbb{R}}^{p}{{\rightarrow}}{\mathbb{R}}^{d}$ are measurable functions such that (2.2) admits a unique strong solution ${\textbf{X}}^{{\textbf{{u}}}}=(X_{i}^{u_{i}})_{i\in[N]}\in\mathcal{S}^{2}({\mathbb{R}}^{dN})$; see (H.1) for the precise conditions. Player $i$ determines their optimal strategy by minimizing the following objective function $J_{i}:\mathcal{A}\to{\mathbb{R}}$:
$\inf_{u_{i}\in\mathcal{A}_{i}}J_{i}(\boldsymbol{u}),\quad\textnormal{with}\quad J_{i}(\boldsymbol{u})\coloneqq\mathbb{E}\left[\int_{0}^{T}f_{i}\left(t,\mathbf{X}_{t}^{\boldsymbol{u}},\boldsymbol{u}_{t}\right)\mathrm{d}t+g_{i}\left(\mathbf{X}_{T}^{\boldsymbol{u}}\right)\right],$ (2.3)
where the running cost $f_{i}:[0,T]\times\mathbb{R}^{dN}\times\mathbb{R}^{kN}\rightarrow\mathbb{R}$ and the terminal cost $g_{i}:\mathbb{R}^{dN}\rightarrow\mathbb{R}$ are given measurable functions.
We denote by $\mathcal{G}$ the game defined by (2.2)–(2.3), and refer to it as a distributed game, as each player’s state is governed solely by their own control. The game $\mathcal{G}$ includes as a special case the game-theoretic models for crowd motion dynamics that will be analyzed in detail in Section 6. In these models, the state process represents each player’s position and/or velocity, and the cost function captures each player’s target region, energy expenditure for traveling, and preferred route, which depends on the spatial distribution of the population. See Section 6 for more details.
Remark 2.1.
Note the distinction between distributed games and distributed controls, the latter of which typically assumes that all players’ states are independent and that each player’s control depends only on their own state (see, e.g., [20]). In contrast, in distributed games players’ states and control processes can be correlated due to shared sources of randomness, such as correlated initial states and common components in the Brownian motions or the Poisson random measures
Remark 2.2.
Note that (2.2) can accommodate linear dependence on the state variable in the drift via a simple change of variables. Indeed, suppose that for each ${\textbf{{u}}}\in\mathcal{A}$, player $i$’s state dynamics $\tilde{X}_{i}^{u_{i}}$ satisfies for all $t\in[0,T]$,
${\mathrm{d}}\tilde{X}_{i,t}=\left(\tilde{a}_{i}(t)\,\tilde{X}_{i,t}+\tilde{b}_{i}(t)\,u_{i,t}\right){\mathrm{d}}t+\tilde{\sigma}_{i}(t)\,{\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}\tilde{\gamma}_{ij}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad\tilde{X}_{i,0}^{u_{i}}=\xi_{i},$ (2.4)
where $\tilde{a}_{i},\tilde{b}_{i},\tilde{\sigma}_{i}$ and $\tilde{\gamma}_{ij}$ are given measurable functions. Then by considering
$X_{i,t}^{u_{i}}\coloneqq e^{-\int_{0}^{t}\tilde{a}_{i}(s)\,{\mathrm{d}}s}\,\tilde{X}_{i,t}^{u_{i}},\quad t\in[0,T],$
the state dynamics (2.4) can be transformed into the simpler form given in (2.2), with the state coefficients and cost functions adjusted by certain deterministic factors. A special case of (2.4) is the following controlled kinetic equation (see e.g, [33, 39]): for all $t\in[0,T]$,
[…]
where $x_{i,t}$ and $v_{i,t}$ denote player $i$’s position and velocity at time $t$, respectively.
Throughout this paper, we impose the following standing regularity condition on the coefficients of (2.2)-(2.3).
H. 1.
For all $i,j\in[N]$, $A_{i}\subset{\mathbb{R}}^{k}$ is convex and $0\in A_{i}$, and $b_{i},\sigma_{i},\gamma_{ij},f_{i}$ and $g_{i}$ are measurable functions satisfying the following conditions:
  1. (1)
    $b_{i}$ and $\sigma_{i}$ are square integrable, and $\sup_{(t,z)\in[0,T]\times{\mathbb{R}}^{p}_{0}}{|\gamma_{ij}(t,z)|}/{\min(1,|z|)}<\infty$.
  2. (2)
[…]
Under Assumption (H.1), for each ${\textbf{{u}}}\in\mathcal{H}^{2}({\mathbb{R}}^{kN})$, (2.2) admits a unique strong solution ${\textbf{X}}^{{\textbf{{u}}}}\in\mathcal{S}^{2}({\mathbb{R}}^{dN})$ (see [26, Theorem 3.1]), and (2.3) is well defined. For ease of exposition, we assume that the action set contains 0, but similar analyses can be extended to a non-empty convex action set (see e.g., [17, 18]).
2.2. NEs and $\alpha$-potential function
We aim to characterize the rational behavior of the players in the distributed game $\mathcal{G}$. To this end, we first recall the notion of an $\varepsilon$-Nash equilibrium, defined as a joint control profile in which no player can improve their performance by more than $\varepsilon$ through any unilateral deviation. The precise definition is given below.
Definition 2.3.
For any $\varepsilon\geq 0$, a control profile $\bar{\textbf{{u}}}=(\bar{u}_{i})_{i\in[N]}\in\mathcal{A}$ is an $\varepsilon$-Nash equilibrium of the game $\mathcal{G}$ if $J_{i}\left(\bar{\textbf{{u}}}\right)\leq J_{i}\left(\left(u_{i},\bar{u}_{-i}\right)\right)+\varepsilon$, for all $i\in[N],u_{i}\in\mathcal{A}_{i}$.
To analyze and compute an approximate NE of the game $\mathcal{G}$, we employ the $\alpha$-potential game framework introduced in [17].
Definition 2.4.
Consider the game $\mathcal{G}$ in (2.2)-(2.3). We say $\mathcal{G}$ is an $\alpha$-potential game for $\alpha\geq 0$ if there exists a function $\Phi:\mathcal{A}\to\mathbb{R}$ such that for all $i\in[N]$, $u_{i},u_{i}^{\prime}\in\mathcal{A}_{i}$ and $u_{-i}\in\mathcal{A}_{-i}$,
$|J_{i}\left(\left(u_{i}^{\prime},u_{-i}\right)\right)-J_{i}\left(\left(u_{i},u_{-i}\right)\right)-\left(\Phi\left(\left(u_{i}^{\prime},u_{-i}\right)\right)-\Phi\left(\left(u_{i},u_{-i}\right)\right)\right)|\leq\alpha.$ (2.5)
Such a function $\Phi$ is called an $\alpha$-potential function for $\mathcal{G}$. In the case where $\alpha=0$, we simply call the game $\mathcal{G}$ a potential game and $\Phi$ a potential function for $\mathcal{G}$.
The main advantage of this framework is that, once such an $\alpha$-potential function $\Phi$ is constructed, finding approximate NEs reduces to solving a single optimization problem: minimizing $\Phi$ over $\mathcal{A}$. This connection is made precise in the following lemma.
Lemma 2.1 ([17, Proposition 2.1]).
Let $\Phi:\mathcal{A}\to{\mathbb{R}}$ be an $\alpha$-potential function of the game $\mathcal{G}$. For each $\varepsilon\geq 0$, if $\bar{\textbf{{u}}}\in\mathcal{A}$ satisfies $\Phi(\bar{\textbf{{u}}})\leq\inf_{{\textbf{{u}}}\in\mathcal{A}}\Phi({\textbf{{u}}})+\varepsilon$, then $\bar{\textbf{{u}}}$ is an $(\alpha+\varepsilon)$-NE of the game $\mathcal{G}$.
As shown in [17], one can construct an $\alpha$-potential function for a stochastic differential game using the linear derivatives of each player’s objective function. For each $i,j\in[N]$, we say $f:\mathcal{A}\to{\mathbb{R}}$ has a linear derivative in $\mathcal{A}_{j}$ if there exists a function $\frac{\delta f}{\delta u_{j}}:\mathcal{H}^{2}({\mathbb{R}}^{kN})\times\mathcal{H}^{2}({\mathbb{R}}^{k})\to{\mathbb{R}}$ such that for all ${\textbf{{u}}}\in\mathcal{A}$, $\frac{\delta f}{\delta u_{j}}({\textbf{{u}}};\cdot)$ is linear and
$\lim_{\varepsilon\searrow 0}\frac{f\left(\left(u_{j}+\varepsilon\left(u_{j}^{\prime}-u_{j}\right),u_{-j}\right)\right)-f({\textbf{{u}}})}{\varepsilon}=\frac{\delta f}{\delta u_{j}}\left({\textbf{{u}}};u_{j}^{\prime}-u_{j}\right),\quad\forall u_{j}^{\prime}\in\mathcal{A}_{j}.$
Similarly, we say $f$ has a second-order linear derivative in $\mathcal{A}_{i}\times\mathcal{A}_{j}$ if $f$ has a linear derivative $\frac{\delta f}{\delta u_{i}}$ in $\mathcal{A}_{i}$, and there exists a function $\frac{\delta^{2}f}{\delta u_{i}\delta u_{j}}:\mathcal{H}^{2}({\mathbb{R}}^{kN})\times\mathcal{H}^{2}({\mathbb{R}}^{k})\times\mathcal{H}^{2}({\mathbb{R}}^{k})\to{\mathbb{R}}$ such that for all ${\textbf{{u}}}\in\mathcal{A}$, $\frac{\delta^{2}f}{\delta u_{i}\delta u_{j}}({\textbf{{u}}};\cdot,\cdot)$ is bilinear and for all $u^{\prime}_{i}\in\mathcal{H}^{2}({\mathbb{R}}^{k})$, $\frac{\delta^{2}f}{\delta u_{i}\delta u_{j}}({\textbf{{u}}};u^{\prime}_{i},\cdot)$ is the linear derivative of ${\textbf{{u}}}\mapsto\frac{\delta f}{\delta u_{i}}({\textbf{{u}}};u^{\prime}_{i})$ in $\mathcal{A}_{j}$.
Using the notion of linear derivatives, the following theorem constructs an $\alpha$-potential function for the game $\mathcal{G}$ and quantify the associate $\alpha$.
Proposition 2.1.
Suppose that for all $i,j\in[N]$, $J_{i}$ has a linear derivative $\frac{\delta J_{i}}{\delta u_{i}}$ in $\mathcal{A}_{i}$, and a second-order linear derivative $\frac{\delta^{2}J_{i}}{\delta u_{i}\delta u_{j}}$ in $\mathcal{A}_{i}\times\mathcal{A}_{j}$. Assume further that for all $u^{\prime}_{i}\in\mathcal{A}_{i}$ and $u^{\prime\prime}_{j}\in\mathcal{A}_{j}$, $\mathcal{A}\ni{\textbf{{u}}}\mapsto\frac{\delta^{2}J_{i}}{\delta u_{i}\delta u_{j}}({\textbf{{u}}};u^{\prime}_{i},u^{\prime\prime}_{j})\in{\mathbb{R}}$ is continuous. Define $\Phi:\mathcal{A}\to{\mathbb{R}}$ by
$\Phi({\textbf{{u}}})=\int_{0}^{1}\sum_{j=1}^{N}\frac{\delta J_{j}}{\delta u_{j}}\left(r{\textbf{{u}}};u_{j}\right)\mathrm{d}r.$ (2.6)
Then $\Phi$ is an $\alpha$-potential function of the game $\mathcal{G}$ with
$\alpha\leq\frac{1}{2}\sup_{i\in[N],u^{\prime}_{i}\in\mathcal{A}_{i},u^{\prime}_{i}\in\mathcal{A}_{j},{\textbf{{u}}}\in\mathcal{A}}\sum_{j=1}^{N}\left|\frac{\delta^{2}J_{i}}{\delta u_{i}\delta u_{j}}\left({\textbf{{u}}};u_{i}^{\prime},u^{\prime\prime}_{j}\right)-\frac{\delta^{2}J_{j}}{\delta u_{j}\delta u_{i}}\left({\textbf{{u}}};u^{\prime\prime}_{j},u_{i}^{\prime}\right)\right|.$ (2.7)
[…]
3. $\alpha$-Potential Function for Distributed Games
This section presents more explicit expressions of the $\alpha$-potential function (2.6) and the corresponding $\alpha$ from Proposition 2.1, expressed in terms of the model coefficients.
The following lemma analytically characterizes the linear derivatives of all players’ objective functions. An important tool is the derivative of each player’s controlled state with respect to her own control, defined by (3.2).
Lemma 3.1.
Suppose (H.1) holds. For all $i\in[N]$, $J_{i}$ has a linear derivative $\frac{\delta J_{i}}{\delta u_{i}}:\mathcal{H}^{2}({\mathbb{R}}^{kN})\times\mathcal{H}^{2}({\mathbb{R}}^{k})\to{\mathbb{R}}$ in $\mathcal{A}_{i}$ satisfying for all ${\textbf{{u}}}\in\mathcal{A}$ and $u^{\prime}_{i}\in\mathcal{A}_{i}$,
$\frac{\delta J_{i}}{\delta u_{i}}\left({\textbf{{u}}};u_{i}^{\prime}\right)$ $=\mathbb{E}\Bigg{[}\int_{0}^{T}\begin{pmatrix}Y^{u^{\prime}_{i}}_{i,t}\\ u_{i,t}^{\prime}\end{pmatrix}^{\top}\begin{pmatrix}\partial_{x_{i}}f_{i}\\ \partial_{a_{i}}f_{i}\end{pmatrix}(t,{\textbf{X}}_{t}^{{\textbf{{u}}}},{\textbf{{u}}}_{t})\,\mathrm{d}t+(Y^{u^{\prime}_{i}}_{i,T})^{\top}(\partial_{x_{i}}g_{i})({\textbf{X}}_{T}^{{\textbf{{u}}}})\Bigg{]},$ (3.1)
where ${\textbf{X}}^{{\textbf{{u}}}}\in\mathcal{S}^{2}({\mathbb{R}}^{dN})$ satisfies (2.2), and $Y_{i}^{u^{\prime}_{i}}\in\mathcal{S}^{2}({\mathbb{R}}^{d})$ satisfies the dynamics
${\mathrm{d}}Y_{i,t}=b_{i}(t)u_{i,t}^{\prime}{\mathrm{d}}t,\quad t\in(0,T];\quad Y_{i,0}=0.$ (3.2)
Moreover, for all $i,j\in[N]$ with $i\not=j$, $J_{i}$ has a second-order linear derivative $\frac{\delta^{2}J_{i}}{\delta u_{i}\delta u_{j}}:\mathcal{H}^{2}({\mathbb{R}}^{kN})\times\mathcal{H}^{2}({\mathbb{R}}^{k})\times\mathcal{H}^{2}({\mathbb{R}}^{k})\to{\mathbb{R}}$ in $\mathcal{A}_{i}\times\mathcal{A}_{j}$ satisfying for all ${\textbf{{u}}}\in\mathcal{A}$, $u^{\prime}_{i}\in\mathcal{A}_{i}$ and $u^{\prime\prime}_{j}\in\mathcal{A}_{j}$,
$\frac{\delta^{2}J_{i}}{\delta u_{i}\delta u_{j}}({\textbf{{u}}};u^{\prime}_{i},u^{\prime\prime}_{j})$ $={\mathbb{E}}\left[\int_{0}^{T}\begin{pmatrix}Y^{u_{i}^{\prime}}_{i,t}\\ u_{i,t}^{\prime}\end{pmatrix}^{\top}\begin{pmatrix}\partial^{2}_{x_{i}x_{j}}f_{i}&\partial^{2}_{x_{i}a_{j}}f_{i}\\ \partial^{2}_{a_{i}x_{j}}f_{i}&\partial^{2}_{a_{i}a_{j}}f_{i}\end{pmatrix}(t,{\textbf{X}}_{t}^{{\textbf{{u}}}},{\textbf{{u}}}_{t})\begin{pmatrix}Y^{u_{j}^{\prime\prime}}_{j,t}\\ u^{\prime\prime}_{j,t}\end{pmatrix}{\mathrm{d}}t\right]$ (3.3)
$\quad+{\mathbb{E}}\left[(Y^{u_{i}^{\prime}}_{i,T})^{\top}(\partial^{2}_{x_{i}x_{j}}g_{i})({\textbf{X}}^{{\textbf{{u}}}}_{T})Y^{u_{j}^{\prime\prime}}_{j,T}\right].$
Lemma 3.1 follows directly from [6, Lemma 4.8], the convexity of $\mathcal{A}_{i}$, and the linearity of the state dynamics (2.2). The expression (3.3) of the second-order derivative $\frac{\delta^{2}J_{i}}{\delta u_{i}\delta u_{j}}$ is simpler than the formula given in [17, Equation 4.6] for general stochastic differential games, due to the fact that player $j$’s control does not affect player $i$’s state evolution.
Leveraging Lemma 3.1, the $\alpha$-potential function given in (2.6) can be expressed as
$\Phi({\textbf{{u}}})$ $=\int_{0}^{1}\sum_{i=1}^{N}\frac{\delta J_{i}}{\delta u_{i}}\left(r{\textbf{{u}}};u_{i}\right){\mathrm{d}}r$ (3.4)
$=\int_{0}^{1}\sum_{i=1}^{N}\mathbb{E}\Bigg{[}\int_{0}^{T}\begin{pmatrix}Y^{u_{i}}_{i,t}\\ u_{i,t}\end{pmatrix}^{\top}\begin{pmatrix}\partial_{x_{i}}f_{i}\\ \partial_{a_{i}}f_{i}\end{pmatrix}(t,{\textbf{X}}_{t}^{r{\textbf{{u}}}},r{\textbf{{u}}}_{t})\,\mathrm{d}t+(Y^{u_{i}}_{i,T})^{\top}(\partial_{x_{i}}g_{i})({\textbf{X}}_{T}^{r{\textbf{{u}}}})\Bigg{]}{\mathrm{d}}r,$
which depends on the aggregated behaviour of ${\textbf{Y}}^{{\textbf{{u}}}}=(Y^{u_{i}}_{i})_{i\in[N]}$ and the family of state processes $({\textbf{X}}^{r{\textbf{{u}}}})_{r\in[0,1]}$ parameterized by $r$. To further simplify the expression (3.4), the following lemma exploits the structure of the state dynamics (2.2) and (3.2), and decomposes ${\textbf{X}}^{r{\textbf{{u}}}}$ into ${\textbf{X}}^{{\textbf{{u}}}}$ and ${\textbf{Y}}^{{\textbf{{u}}}}$.
Lemma 3.2.
Suppose (H.1) holds. For all ${\textbf{{u}}}\in\mathcal{H}^{2}({\mathbb{R}}^{kN})$ and $r\in[0,1]$, ${\textbf{X}}^{r{\textbf{{u}}}}={\textbf{X}}^{{\textbf{{u}}}}-(1-r){\textbf{Y}}^{{\textbf{{u}}}}$.
The proof simply follows by noting that the process $\tilde{{\textbf{X}}}\coloneqq{\textbf{X}}^{{\textbf{{u}}}}-(1-r){\textbf{Y}}^{{\textbf{{u}}}}$ has the same initial condition and satisfies the same dynamics as ${{\textbf{X}}}^{r{\textbf{{u}}}}$.
Based on Lemma 3.2, the following theorem simplifies the expression (3.4) of the $\alpha$-potential function, and derives an explicit upper bound for $\alpha$ in terms of the model coefficients.
Theorem 3.1.
Suppose (H.1) holds. The function $\Phi:\mathcal{A}^{(N)}\to{\mathbb{R}}$ in (2.6) can be expressed as
$\Phi(\boldsymbol{u})=\mathbb{E}$ $\left[\int_{0}^{T}F(t,{\textbf{X}}^{{\textbf{{u}}}}_{t},{\textbf{Y}}^{{\textbf{{u}}}}_{t},{\textbf{{u}}}_{t}){\mathrm{d}}t+G({\textbf{X}}^{{\textbf{{u}}}}_{T},{\textbf{Y}}^{{\textbf{{u}}}}_{T})\right],$ (3.5)
where for each ${\textbf{{u}}}=(u_{i})_{i\in[N]}\in\mathcal{A}$, ${\textbf{X}}^{{\textbf{{u}}}}=(X^{u_{i}}_{i})_{i\in[N]}$ and ${\textbf{Y}}^{{\textbf{{u}}}}=(Y^{u_{i}}_{i})_{i\in[N]}$ satisfy (2.2) and (3.2), respectively, and $F:[0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{kN}\to{\mathbb{R}}$ and $G:{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}\to{\mathbb{R}}$ satisfy for all $t\in[0,T]$, $x=(x_{i})_{i\in[N]},y=(y_{i})_{i\in[N]}\in{\mathbb{R}}^{dN}$ and $a=(a_{i})_{i\in[N]}\in{\mathbb{R}}^{kN}$,
$\begin{split}F(t,x,y,a)&\coloneqq\sum_{i=1}^{N}\int_{0}^{1}\binom{y_{i}}{a_{i}}^{\top}\binom{\partial_{x_{i}}f_{i}}{\partial_{a_{i}}f_{i}}\left(t,x-(1-r)y,ra\right){\mathrm{d}}r,\\ G(x,y)&\coloneqq\sum_{i=1}^{N}\int_{0}^{1}y_{i}^{\top}\left(\partial_{x_{i}}g_{i}\right)\left(x-(1-r)y\right){\mathrm{d}}r.\end{split}$ (3.6)
Moreover, $\Phi$ is an $\alpha$-potential function of the game $\mathcal{G}$ with
$\alpha$ $\leq\frac{1}{2}\sup_{i\in[N]}\sum_{j\in[N]\setminus\{i\}}U_{i}U_{j}\Bigg{(}TB_{i}B_{j}\|\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}+T^{\frac{1}{2}}B_{i}\|\partial^{2}_{x_{i}a_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}+T^{\frac{1}{2}}B_{j}\|\partial^{2}_{a_{i}x_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}$ (3.7)
$\qquad+\|\partial^{2}_{a_{i}a_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}+B_{i}B_{j}\|\partial^{2}_{x_{i}x_{j}}\Delta^{g}_{i,j}\|_{L^{\infty}}\Bigg{)},$
where for all $i,j\in[N]$ with $i\not=j$, $\Delta^{f}_{i,j}\coloneqq f_{i}-f_{j}$, $\Delta^{g}_{i,j}\coloneqq g_{i}-g_{j}$, $B_{i}\coloneqq\|b_{i}\|_{L^{2}}$ and $U_{i}\coloneqq\sup_{u_{i}\in\mathcal{A}_{i}}\|u_{i}\|_{\mathcal{H}^{2}}$.
Proof.
The expression (3.5) follows by substituting the expression ${\textbf{X}}^{r{\textbf{{u}}}}={\textbf{X}}^{{\textbf{{u}}}}-(1-r){\textbf{Y}}^{{\textbf{{u}}}}$ into Lemma 3.2, and applying Fubini’s theorem.
To get an upper bound of $\alpha$, by Lemma 3.1,
$\begin{split}&\frac{\delta^{2}J_{i}}{\delta u_{i}\delta u_{j}}({\textbf{{u}}};u^{\prime}_{i},u^{\prime\prime}_{j})-\frac{\delta^{2}J_{j}}{\delta u_{j}\delta u_{i}}({\textbf{{u}}};u^{\prime\prime}_{j},u^{\prime}_{i})\\ &={\mathbb{E}}\left[\int_{0}^{T}\begin{pmatrix}Y^{u_{i}^{\prime}}_{i,t}\\ u_{i,t}^{\prime}\end{pmatrix}^{\top}\begin{pmatrix}\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j}&\partial^{2}_{x_{i}a_{j}}\Delta^{f}_{i,j}\\ \partial^{2}_{a_{i}x_{j}}\Delta^{f}_{i,j}&\partial^{2}_{a_{i}a_{j}}\Delta^{f}_{i,j}\end{pmatrix}(t,\cdot)\begin{pmatrix}Y^{u_{j}^{\prime\prime}}_{j,t}\\ u^{\prime\prime}_{j,t}\end{pmatrix}{\mathrm{d}}t+(Y^{u_{i}^{\prime}}_{i,T})^{\top}(\partial^{2}_{x_{i}x_{j}}\Delta^{g}_{i,j})({\textbf{X}}^{{\textbf{{u}}}}_{T})Y^{u_{j}^{\prime\prime}}_{j,T}\right],\end{split}$ (3.8)
where we write $\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j}(t,\cdot)=\partial^{2}_{x_{i}x_{j}}(f_{i}-f_{j})(t,{\textbf{X}}_{t}^{\textbf{{u}}},{\textbf{{u}}}_{t})$ and similarly for other derivatives. Moreover, by (3.2), for any $t\in[0,T]$, $Y_{i,t}^{u_{i}^{\prime}}=\int_{0}^{t}b_{i}(v)u_{i}^{\prime}(v){\mathrm{d}}v$, and hence by the Cauchy-Schwarz inequality,
$\mathbb{E}\left[\left|Y_{i,t}^{u_{i}^{\prime}}\right|^{2}\right]$ $=\mathbb{E}\left[\left|\int_{0}^{t}b_{i}(v)u_{i}^{\prime}(v){\mathrm{d}}v\right|^{2}\right]\leq\mathbb{E}\left[\int_{0}^{t}|b_{i}(v)|^{2}{\mathrm{d}}v\int_{0}^{t}|u_{i}^{\prime}(v)|^{2}{\mathrm{d}}v\right]=\|b_{i}\|_{L^{2}}^{2}\|u_{i}^{\prime}\|_{\mathcal{H}^{2}}^{2}.$ (3.9)
Thus $\|Y_{i}^{u_{i}^{\prime}}\|^{2}_{\mathcal{H}^{2}}\leq T\sup_{t\in[0,T]}\mathbb{E}\left[\left|Y_{i,t}^{u_{i}^{\prime}}\right|^{2}\right]\leq T\|b_{i}\|_{L^{2}}^{2}\|u_{i}^{\prime}\|_{\mathcal{H}^{2}}^{2}$.
We now estimate each term in (3.8). Observe that for all $t\in[0,T]$,
$\begin{split}&\left|{\mathbb{E}}\left[\int_{0}^{T}\left(Y^{u^{\prime}_{i}}_{i,t}\right)^{\top}(\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j})(t,\cdot)Y^{u^{\prime\prime}_{j}}_{j,t}{\mathrm{d}}t\right]\right|\leq{\mathbb{E}}\left[\int_{0}^{T}|Y^{u^{\prime}_{i}}_{i,t}|\|(\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j})(t,\cdot)\|_{\textrm{sp}}|Y^{u^{\prime\prime}_{j}}_{j,t}|{\mathrm{d}}t\right]\\ &\leq\|\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}{\mathbb{E}}\left[\int_{0}^{T}|Y^{u^{\prime}_{i}}_{i,t}||Y^{u^{\prime\prime}_{j}}_{j,t}|{\mathrm{d}}t\right]\leq\|\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}\|Y^{u^{\prime}_{i}}_{i}\|_{\mathcal{H}^{2}}\|Y^{u^{\prime\prime}_{j}}_{j}\|_{\mathcal{H}^{2}}\\ &\leq T\|\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}\|b_{i}\|_{L^{2}}\|b_{j}\|_{L^{2}}\|u_{i}^{\prime}\|_{\mathcal{H}^{2}}\|u_{j}^{\prime\prime}\|_{\mathcal{H}^{2}}.\end{split}$ (3.10)
Similarly, we have
$\begin{split}&\Bigg{|}{\mathbb{E}}\Bigg{[}\int_{0}^{T}\left(Y^{u^{\prime}_{i}}_{i,t}\right)^{\top}(\partial^{2}_{x_{i}a_{j}}\Delta^{f}_{i,j})(t,\cdot)u_{j,t}^{\prime\prime}{\mathrm{d}}t\Bigg{]}\Bigg{|}\leq\|\partial^{2}_{x_{i}a_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}\|Y^{u^{\prime}_{i}}_{i}\|_{\mathcal{H}^{2}}\|u_{j}^{\prime\prime}\|_{\mathcal{H}^{2}}\\ &\leq T^{\frac{1}{2}}\|\partial^{2}_{x_{i}a_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}\|b_{i}\|_{L^{2}}\|u_{i}^{\prime}\|_{\mathcal{H}^{2}}\|u_{j}^{\prime\prime}\|_{\mathcal{H}^{2}},\end{split}$ (3.11)
and that
$\begin{split}\left|{\mathbb{E}}\left[\int_{0}^{T}\left(u^{\prime}_{i,t}\right)^{\top}(\partial^{2}_{a_{i}x_{j}}\Delta^{f}_{i,j})(t,\cdot)Y_{j,t}^{u^{\prime\prime}_{j}}{\mathrm{d}}t\right]\right|&\leq T^{\frac{1}{2}}\|\partial^{2}_{a_{i}x_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}\|b_{j}\|_{L^{2}}\|u_{i}^{\prime}\|_{\mathcal{H}^{2}}\|u_{j}^{\prime\prime}\|_{\mathcal{H}^{2}},\\ \left|{\mathbb{E}}\left[\int_{0}^{T}\left(u^{\prime}_{i,t}\right)^{\top}(\partial^{2}_{a_{i}a_{j}}\Delta^{f}_{i,j})(t,\cdot)u_{j,t}^{\prime\prime}{\mathrm{d}}t\right]\right|&\leq\|\partial^{2}_{a_{i}a_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}\|u_{i}^{\prime}\|_{\mathcal{H}^{2}}\|u_{j}^{\prime\prime}\|_{\mathcal{H}^{2}}.\end{split}$ (3.12)
Finally, we have
${\mathbb{E}}\left[(Y^{u_{i}^{\prime}}_{i,T})^{\top}(\partial^{2}_{x_{i}x_{j}}\Delta^{g}_{i,j})({\textbf{X}}^{{\textbf{{u}}}}_{T})Y^{u_{j}^{\prime\prime}}_{j,T}\right]$ $\leq\|\partial^{2}_{x_{i}x_{j}}\Delta^{g}_{i,j}\|_{L^{\infty}}{\mathbb{E}}\left[|Y^{u_{i}^{\prime}}_{i,T}||Y^{u_{j}^{\prime\prime}}_{j,T}|\right]$ (3.13)
$\leq\|\partial^{2}_{x_{i}x_{j}}\Delta^{g}_{i,j}\|_{L^{\infty}}\|b_{i}\|_{L^{2}}\|b_{j}\|_{L^{2}}\|u_{i}^{\prime}\|_{\mathcal{H}^{2}}\|u_{j}^{\prime\prime}\|_{\mathcal{H}^{2}}.$
Combining (3.10), (3.11), (3.12), (3.13), and Proposition 2.1 yields the desired result. ∎
Compared with (3.4), (3.5) isolates the contribution of $r$ and expresses the $\alpha$-potential function only in terms of ${\textbf{X}}^{\textbf{{u}}}$ and ${\textbf{Y}}^{\textbf{{u}}}$. This reformulation enables the use of standard control techniques to minimize the $\alpha$-potential function, thereby simplifying the computation of approximate Nash equilibria for the game $\mathcal{G}$.
To see it, recall that the objective function (3.4) depends on the aggregated behavior of the state processes ${\textbf{X}}^{r{\textbf{{u}}}}$ with respect to $r\in[0,1]$. This parameter $r$ acts as a uniformly distributed noise independent of ${\mathbb{F}}$. As shown in [17], to find a minimizer of (3.4) that is adapted to ${\mathbb{F}}$, one must lift the problem into a conditional McKean–Vlasov control framework, where the state variable becomes the conditional law $\mathcal{L}({\textbf{X}}^{r{\textbf{{u}}}},{\textbf{Y}}^{{\textbf{{u}}}},r\mid{\mathbb{F}})$. The resulting optimal control is characterized by an infinite-dimensional Hamilton–Jacobi–Bellman (HJB) equation defined on the space of probability measures.
In contrast, the reformulated objective (3.5) depends only on the $2dN$-dimensional state variables $({\textbf{X}}^{{\textbf{{u}}}},{\textbf{Y}}^{{\textbf{{u}}}})$. The corresponding optimal control can then be characterized by a standard HJB equation defined on the space ${\mathbb{R}}^{2dN}$, as will be shown in Section 4.
We further remark that when the upper bound in (3.7) is zero, the game $\mathcal{G}$ becomes a potential game, and its Nash equilibria can be obtained by minimizing a potential function that involves only on the state variable ${\textbf{X}}^{{\textbf{{u}}}}$, as defined in (3.15).
Theorem 3.2.
Suppose (H.1) holds, and for all $i,j\in[N]$ with $i\not=j$,
$\partial^{2}_{x_{i}x_{j}}f_{i}=\partial^{2}_{x_{i}x_{j}}f_{j},\quad\partial^{2}_{a_{i}x_{j}}f_{i}=\partial^{2}_{a_{i}x_{j}}f_{j},\quad\partial^{2}_{a_{i}a_{j}}f_{i}=\partial^{2}_{a_{i}a_{j}}f_{j},\quad\partial^{2}_{x_{i}x_{j}}g_{i}=\partial^{2}_{x_{i}x_{j}}g_{j}.$ (3.14)
Then the game $\mathcal{G}$ is a potential game with a potential function defined by
$\bar{\Phi}({\textbf{{u}}})\coloneqq\mathbb{E}\left[\int_{0}^{T}\bar{F}(t,{\textbf{X}}^{{\textbf{{u}}}}_{t},{\textbf{{u}}}_{t}){\mathrm{d}}t+\bar{G}({\textbf{X}}^{{\textbf{{u}}}}_{T})\right],$ (3.15)
where ${\textbf{X}}^{{\textbf{{u}}}}$ satisfies (2.2), $\bar{F}:[0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{kN}\to{\mathbb{R}}$ and $\bar{G}:{\mathbb{R}}^{dN}\to{\mathbb{R}}$ satisfy for all $t\in[0,T]$, $x=(x_{i})_{i\in[N]}\in{\mathbb{R}}^{dN}$ and $a=(a_{i})_{i\in[N]}\in{\mathbb{R}}^{kN}$,
$\begin{split}\bar{F}(t,x,a)&\coloneqq\sum_{i=1}^{N}\int_{0}^{1}\binom{x_{i}}{a_{i}}^{\top}\binom{\partial_{x_{i}}f_{i}}{\partial_{a_{i}}f_{i}}\left(t,rx,ra\right){\mathrm{d}}r,\quad G(x)\coloneqq\sum_{i=1}^{N}\int_{0}^{1}x_{i}^{\top}\left(\partial_{x_{i}}g_{i}\right)\left(rx\right){\mathrm{d}}r.\end{split}$
Moreover, for all ${\textbf{{u}}}\in\mathcal{A}$,
$\bar{\Phi}({\textbf{{u}}})=\Phi({\textbf{{u}}})+\mathbb{E}\left[\int_{0}^{T}\bar{F}(t,{\textbf{X}}^{\boldsymbol{0}}_{t},\boldsymbol{0}){\mathrm{d}}t+\bar{G}({\textbf{X}}^{\boldsymbol{0}}_{T})\right],$ (3.16)
where $\Phi$ is defined in (3.5) and $\boldsymbol{0}$ is the constant process taking the value zero at all times.
Proof.
Under the symmetry condition (3.14), the fact that the function $\bar{\Phi}$ in (3.15) is a potential function for the game $\mathcal{G}$ follows from analogous arguments to those used for distributed games with Markov policies in [18, Theorem 3.2]. Since both $\bar{\Phi}$ and $\Phi$ are potential functions for the game $\mathcal{G}$, it holds for all $u\in\mathcal{A}$,
$\bar{\Phi}({\textbf{{u}}})-\Phi({\textbf{{u}}})=\bar{\Phi}(\boldsymbol{0})-\Phi(\boldsymbol{0}).$
To see it, assume without loss of generality that $N=2$. Then for all ${\textbf{{u}}}=(u_{i})_{i=1}^{2}\in\mathcal{A}$, using the definition (2.5) of a potential function,
$\bar{\Phi}({\textbf{{u}}})-\Phi({\textbf{{u}}})$ $=\bar{\Phi}((u_{1},u_{2}))-\bar{\Phi}((0,u_{2}))+\bar{\Phi}((0,u_{2}))-\bar{\Phi}(\boldsymbol{0})+\bar{\Phi}(\boldsymbol{0})$
$\quad-\left(\Phi((u_{1},u_{2}))-\Phi((0,u_{2}))+\Phi((0,u_{2}))-\Phi(\boldsymbol{0})+\Phi(\boldsymbol{0})\right)$
$=J_{1}((u_{1},u_{2}))-J_{1}((0,u_{2}))+J_{2}((0,u_{2}))-J_{2}((0,0))+\bar{\Phi}(\boldsymbol{0})$
$\quad-\left(J_{1}((u_{1},u_{2}))-J_{1}((0,u_{2}))+J_{2}((0,u_{2}))-J_{2}((0,0))-J_{2}(0,0)+\Phi(\boldsymbol{0})\right)$
$=\bar{\Phi}(\boldsymbol{0})-\Phi(\boldsymbol{0}).$
The desired identity (3.16) then follows from the fact that ${\textbf{Y}}^{\boldsymbol{0}}_{t}=0$ for all $t\in[0,T]$, and $F(t,x,0,0)=G(x,0)$ for all $(t,x)\in[0,T]\times{\mathbb{R}}^{dN}$. ∎
4. Optimize $\alpha$-Potential Function for $\alpha$-NE
Given the $\alpha$-potential function $\Phi$ defined in (3.5), this section characterizes its minimizer over the admissible control space $\mathcal{A}$, which in turn constructs analytically an $\alpha$-NE for the distributed game $\mathcal{G}$. We adopt a dynamic programming approach that characterizes the minimizer of the $\alpha$-potential function in feedback form via solutions to suitable HJB integro-partial differential equations. This characterization offers a theoretical foundation for developing policy gradient algorithms to solve the distributed game $\mathcal{G}$; see Section 5 for details.
More precisely, we consider the following control problem
$\inf_{{\textbf{{u}}}\in\mathcal{A}}\Phi({\textbf{{u}}}),\quad\Phi({\textbf{{u}}})=\mathbb{E}$ $\left[\int_{0}^{T}F(t,{\textbf{X}}^{{\textbf{{u}}}}_{t},{\textbf{Y}}^{{\textbf{{u}}}}_{t},{\textbf{{u}}}_{t}){\mathrm{d}}t+G({\textbf{X}}^{{\textbf{{u}}}}_{T},{\textbf{Y}}^{{\textbf{{u}}}}_{T})\right],$ (4.1)
where $\mathcal{A}$ is the set of admissible controls given by
$\mathcal{A}\coloneqq\{u:\Omega\times[0,T]\to A\mid u\in\mathcal{H}^{2}({\mathbb{R}}^{kN})\},$
$F$ and $G$ are defined in (3.6), and $({\textbf{X}}^{{\textbf{{u}}}},{\textbf{Y}}^{{\textbf{{u}}}})$ satisfy the following state dynamics:
${\mathrm{d}}{\textbf{X}}_{t}&=b(t){\textbf{{u}}}_{t}{\mathrm{d}}t+\sigma(t){\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}{\gamma}_{j}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad{\textbf{X}}_{0}=\boldsymbol{\xi},\\ {\mathrm{d}}{\textbf{Y}}_{t}&=b(t){\textbf{{u}}}_{t}{\mathrm{d}}t,\quad{\textbf{Y}}_{0}=0,\end{aligned}\right.$ (4.2)
where $\boldsymbol{\xi}=(\xi^{\top}_{1},\ldots,\xi^{\top}_{N})^{\top}$, and for all $t\in[0,T]$ and $z\in{\mathbb{R}}^{p}_{0}$,
$b(t)\coloneqq\operatorname{diag}(b_{1}(t),\cdots,b_{N}(t))\in{\mathbb{R}}^{dN\times kN},\quad\sigma(t)\coloneqq\begin{pmatrix}\sigma_{1}(t)\\ \vdots\\ \sigma_{N}(t)\end{pmatrix}\in{\mathbb{R}}^{dN\times n},\quad\gamma_{j}(t,z)\coloneqq\begin{pmatrix}\gamma_{1j}(t,z)\\ \vdots\\ \gamma_{Nj}(t,z)\end{pmatrix}\in{\mathbb{R}}^{dN}.$
4.1. Verification theorem
The minimizer of (4.1) can be constructed by standard verification results. To see it, let $\mathcal{C}^{1,2,1}([0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN})$ be the space of functions $\phi=\phi(t,x,y):[0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}\to{\mathbb{R}}$ such that $\partial_{t}\phi$, $\partial_{x}\phi$, $\partial^{2}_{xx}\phi$, and $\partial_{y}\phi$ exist and are continuous. For all $a\in A$ and $\phi\in\mathcal{C}^{1,2,1}([0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN})$, define the operator $\mathbb{L}^{a}\phi:[0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}{{\rightarrow}}{\mathbb{R}}$ by
$\mathbb{L}^{a}\phi(t,x,y)\coloneqq$ $\left(b(t)a\right)^{\top}\left(\partial_{x}\phi(t,x,y)+\partial_{y}\phi(t,x,y)\right)+\frac{1}{2}\operatorname{tr}\left(\sigma(t)\sigma(t)^{\top}\partial_{xx}^{2}\phi(t,x,y)\right)$
$+\sum_{j=1}^{m}\int_{{\mathbb{R}}^{p}_{0}}\left(\phi(t,x+\gamma_{j}(t,z),y)-\phi(t,x,y)-\partial_{x}\phi(t,x,y)^{\top}\gamma_{j}(t,z)\right)\nu_{j}({\mathrm{d}}z),$
and define the associated Hamiltonian by
$H(t,x,y,\phi,a)=\mathbb{L}^{a}\phi(t,x,y)+F(t,x,y,a).$
The HJB equation associated with (4.1) is given by
[…]
We now present the verification theorem, which constructs an optimal control of (4.1) (and hence an $\alpha$-NE of the game $\mathcal{G}$) in a feedback form using a smooth solution to the HJB equation (4.3).
Theorem 4.1.
Suppose (H.1) holds. Assume that there exists $v\in C^{1,2,1}([0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN})$ such that $\inf_{a\in A}H(t,x,y,v,a)\in{\mathbb{R}}$ for all $(t,x,y)\in[0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}$, and $v$ satisfies the HJB equation (4.3). Assume further that there exists a measurable map $\hat{a}:[0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}\to A$ such that
$\hat{a}(t,x,y)=\arg\min_{a\in A}H(t,x,y,v,a),\quad(t,x,y)\in[0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN},$ (4.4)
the corresponding controlled dynamics
${\mathrm{d}}{\textbf{X}}_{t}&=b(t)\hat{a}(t,{\textbf{X}}_{t},{\textbf{Y}}_{t}){\mathrm{d}}t+\sigma(t){\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}{\gamma}_{j}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad{\textbf{X}}_{0}=\boldsymbol{\xi},\\ {\mathrm{d}}{\textbf{Y}}_{t}&=b(t)\hat{a}(t,{\textbf{X}}_{t},{\textbf{Y}}_{t}){\mathrm{d}}t,\quad{\textbf{Y}}_{0}=0,\end{aligned}\right.$ (4.5)
has a square integrable strong solution $(\hat{{\textbf{X}}},\hat{{\textbf{Y}}})$ and that the control $\hat{\textbf{{u}}}_{t}\coloneqq\hat{a}(t,\hat{{\textbf{X}}}_{t},\hat{{\textbf{Y}}}_{t})$, $t\in[0,T]$, is in $\mathcal{H}^{2}({\mathbb{R}}^{kN})$. Then $v(0,\xi)=\inf_{{\textbf{{u}}}\in\mathcal{A}}\Phi({\textbf{{u}}})$, and $\hat{\textbf{{u}}}$ is an optimal control of (4.1) and an $\alpha$-NE of the distributed game $\mathcal{G}$, with $\alpha$ given in (3.7).
[…]
The proof of Theorem 4.1 follows from standard verification arguments for classical stochastic control problems (see, e.g., [46, Chapter 5]). The first step is to show that $\Phi({\textbf{{u}}})\geq v(0,\xi,0)$ for all ${\textbf{{u}}}\in\mathcal{A}$, by applying Itô’s formula for jump-diffusion processes to the function $t\mapsto v(t,{\textbf{X}}^{{\textbf{{u}}}}_{t},{\textbf{Y}}^{{\textbf{{u}}}}_{t})$ and using the HJB equation (4.3) satisfied by $v$. The second step is to show that $\Phi(\hat{{\textbf{{u}}}})=v(0,\xi,0)$ due to the definition (4.4) of $\hat{a}$, which implies the optimality of $\hat{\textbf{{u}}}$.
4.2. Viscosity characterization
In the case where the HJB equation (4.3) does not admit a classical solution, we can characterize the value function of (4.1) as the continuous viscosity solution of (4.3). To this end, define the value function starting from time $t\in[0,T]$ and state $(x,y)\in{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}$ by
$V(t,x,y)\coloneqq\inf_{{\textbf{{u}}}\in\mathcal{A}}\mathbb{E}$ $\left[\int_{t}^{T}F(s,{\textbf{X}}^{{\textbf{{u}}}}_{t},{\textbf{Y}}^{{\textbf{{u}}}}_{s},{\textbf{{u}}}_{s}){\mathrm{d}}s+G({\textbf{X}}^{{\textbf{{u}}}}_{T},{\textbf{Y}}^{{\textbf{{u}}}}_{T})\bigg{|}X^{{\textbf{{u}}}}_{t}=x,Y^{{\textbf{{u}}}}_{t}=y\right].$ (4.6)
We impose the following assumptions, which are standard in the literature for establishing the uniqueness of viscosity solutions (see e.g., [35, 12, 21]).
H. 2.
Assume the setting in (H.1). For all $i,j\in[N]$, $A_{i}$ is compact, $b_{i}$ $\sigma_{i}$, and $\gamma_{ij}$ are continuous in $t$, and $\partial_{x_{i}}f_{i}$ and $\partial_{a_{i}}f_{i}$ are continuous in all variables.
Now we identify the value function $V$ defined by (4.6) as the unique viscosity solution to (4.3). As in [21, Definition 2.1], we say a function $v:[0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}\to{\mathbb{R}}$ is a viscosity subsolution (resp. supersolution) of (4.3) if $v$ is upper semicontinuous (resp. lower semicontinuous) and for every $(t_{0},x_{0},y_{0})\in[0,T)\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}$ and $\phi\in C^{1,2,1}([0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN})$ such that $\phi-v$ attains its minimum (resp. maximum) at $(t_{0},x_{0},y_{0})$,
[…]
Theorem 4.2.
Suppose (H.2) holds. The function $V$ defined by (4.6) is the unique viscosity solution of the HJB equation (4.3) in the class of continuous functions with at most quadratic growth in $(x,y)$, in the sense that $V$ is a viscosity sub- and supersolution of (4.3) with terminal condition $V(T,x,y)=G(x,y)$.
Theorem 4.2 is a fairly standard result; however, as most existing references focus on the case where the cost functions $F$ and $G$ are globally Lipschitz continuous (see, e.g., [35, 12, 4]), we include a brief sketch of the proof below for completeness. Under (H.2), the functions $F$ and $G$ defined in (3.6) are continuous and it holds for some $C\geq 0$ that for all $t\in[0,T]$ and $(x,y),(x^{\prime},y^{\prime})\in{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}$, $|F(t,x,y)|+|G(x,y)|\leq C(1+|x|^{2}+|y|^{2})$ and
$|F(t,x,y)-F(t,x^{\prime},y^{\prime})|+|G(x,y)-G(x^{\prime},y^{\prime})|\leq C(1+|x|+|y|+|x^{\prime}|+|y^{\prime}|)(|x-x^{\prime}|+|y-y^{\prime}|).$
[…]
5. Policy Gradient Algorithm for $\alpha$-NE
Theorem 4.1 characterizes an open-loop $\alpha$-NE for the distributed game $\mathcal{G}$ in the feedback form with respect to the state process X and the sensitivity process Y. The feedback controls therein are constructed from solutions to the corresponding HJB equations, which may not admit closed-form expressions.
In this section, we propose a policy gradient algorithm to compute the $\alpha$-NE for the distributed game $\mathcal{G}$. The algorithm searches for the $\alpha$-NE by directly minimizing the $\alpha$-potential function (3.5) over suitable parametric families. For clarity of exposition, we present the algorithm under the assumption that the jump measures $(\nu_{j})_{j=1}^{m}$ in (2.2) are finite, i.e.,
$\nu_{j}({\mathbb{R}}^{p}_{0})<\infty,\quad\forall j=1,\ldots,m.$
Problems involving singular jump measures with infinitely many jumps can be reduced to ones with finite-activity measures by applying the standard diffusion approximation (see, e.g., [10, 13, 38]). This approach involves truncating the singular measures at a given threshold and approximating the small-jump component using a modified diffusion coefficient. The approximation error depends on the choice of truncation threshold and the singularity of the jump measures $(\nu_{i})_{i=1}^{m}$ near zero (see e.g., [13, Lemma C.3]).
The algorithm begins by approximating the NE policy given in Theorem 4.1 using a sufficiently expressive parametric family (e.g., a family of deep neural networks). Specifically, we consider a family of policy profiles $\phi_{\theta}:[0,T]\times\mathbb{R}^{dN}\times\mathbb{R}^{dN}\to A$ with weights $\theta\in\mathbb{R}^{L}$, and consider for each $\theta\in{\mathbb{R}}^{L}$,
$\Phi(\theta)\coloneqq\mathbb{E}$ $\left[\int_{0}^{T}F(t,{\textbf{X}}^{\theta}_{t},{\textbf{Y}}^{\theta}_{t},\phi_{\theta}(t,\mathbf{X}^{\theta}_{t},\mathbf{Y}^{\theta}_{t})){\mathrm{d}}t+G({\textbf{X}}^{\theta}_{T},{\textbf{Y}}^{\theta}_{T})\right],$ (5.1)
where $({\textbf{X}}^{\theta},{\textbf{Y}}^{\theta})$ are the state and sensitivity processes satisfying the following dynamics:
${\mathrm{d}}{\textbf{X}}_{t}&=b(t)\phi_{\theta}(t,{\textbf{X}}_{t},{\textbf{Y}}_{t}){\mathrm{d}}t+\sigma(t){\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}{\gamma}_{j}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad{\textbf{X}}_{0}=\boldsymbol{\xi},\\ {\mathrm{d}}{\textbf{Y}}_{t}&=b(t)\phi_{\theta}(t,{\textbf{X}}_{t},{\textbf{Y}}_{t}){\mathrm{d}}t,\quad{\textbf{Y}}_{0}=0,\end{aligned}\right.$ (5.2)
That is, we restrict the control problem (4.1) on the set of controls ${\textbf{{u}}}_{t}=\phi_{\theta}(t,\mathbf{X}^{\theta}_{t},\mathbf{Y}^{\theta}_{t})$, $t\in[0,T]$, induced by $\phi_{\theta}$.
We seek an optimal policy that minimizes (5.1), which yields an approximate NE of the distributed game $\mathcal{G}$ as shown in Lemma 2.1 and Theorem 4.1. This is achieved by performing gradient descent of (5.1) with respect to the weights $\theta$ based on simulated trajectories of (5.2). More precisely, given a fixed policy $\phi_{\theta}$, we consider the following Euler-Maruyama approximation of (5.2) on the time grid $\pi_{P}\coloneqq\{0=t_{0}<\ldots<t_{P}=T\}$ for some $P\in{\mathbb{N}}$: for all $i\in[N]$, let $X^{\theta}_{i,0}=\xi_{i}$ and $Y^{\theta}_{i,0}=0$, and for all $\ell=0,\ldots,P-1$,
$X^{\theta}_{i,t_{\ell+1}}$ $=X^{\theta}_{i,t_{\ell}}+b_{i}(t_{\ell})\phi_{\theta}(t_{\ell},{\textbf{X}}^{\theta}_{t_{\ell}},{\textbf{Y}}^{\theta}_{t_{\ell}})\Delta_{\ell}+\sigma_{i}(t_{\ell})\Delta W_{\ell}$ (5.3)
$\quad+\sum_{j=1}^{m}\left(\sum_{k=N_{j,\ell}+1}^{N_{j,\ell+1}}\gamma_{ij}(t_{\ell},z_{k})-\Delta_{\ell}\int_{{\mathbb{R}}^{p}_{0}}\gamma_{ij}(t_{\ell},z)\nu({\mathrm{d}}z)\right),$
$Y^{\theta}_{i,t_{\ell+1}}$ $=Y^{\theta}_{i,t_{\ell}}+b_{i}(t_{\ell})\phi_{\theta}(t_{\ell},{\textbf{X}}^{\theta}_{t_{\ell}},{\textbf{Y}}^{\theta}_{t_{\ell}})\Delta_{\ell},\quad{\textbf{X}}^{\theta}_{t_{\ell}}=(X^{\theta}_{i,t_{\ell}})_{i\in[N]},\quad{\textbf{Y}}^{\theta}_{t_{\ell}}=(Y^{\theta}_{i,t_{\ell}})_{i\in[N]},$
where $\Delta_{\ell}\coloneqq t_{\ell+1}-t_{\ell}$, $\Delta W_{\ell}\coloneqq W_{t_{\ell+1}}-W_{t_{\ell}}$, $N_{j,\ell}$ denotes the number of jumps of the $j$-th Poisson random measure occurring over the time interval $[0,t_{\ell}]$, and $z_{k}$ is the size of the $k$-th jump sampled from the distribution $\nu/\nu({\mathbb{R}}^{p}_{0})$. Let $({\textbf{X}}^{\theta,(m)},{\textbf{Y}}^{\theta,(m)})_{m=1}^{M}$, $M\in{\mathbb{N}}$, be independent trajectories of (5.3) with policy $\phi_{\theta}$, and define the following empirical approximation of (5.1)
$\Phi_{M}(\theta)\coloneqq\frac{1}{M}\sum_{m=1}^{M}\left[\sum_{\ell=0}^{P-1}F\left({\textbf{X}}^{\theta,(m)}_{t_{\ell}},{\textbf{Y}}^{\theta,(m)}_{t_{\ell}},\phi_{\theta}\left(t_{\ell},{\textbf{X}}^{\theta,(m)}_{t_{\ell}},{\textbf{Y}}^{\theta,(m)}_{t_{\ell}}\right)\right)\Delta_{\ell}+G\left({\textbf{X}}^{\theta,(m)}_{t_{P}},{\textbf{Y}}^{\theta,(m)}_{t_{P}}\right)\right].$ (5.4)
By choosing a sufficiently large $M$ and minimizing (5.4) over $\theta$, we obtain an approximate minimizer of the $\alpha$-potential function, and consequently an approximate NE for the game $\mathcal{G}$.
Here we summarize the above policy gradient algorithm for the $\alpha$-potential game $\mathcal{G}$. For simplicity, we present a version of the algorithm that minimizes (5.4) using a mini-batch stochastic gradient descent method. In practice, more sophisticated variants of stochastic gradient descent (such as Adam [24]) can be employed to optimize (5.4) more efficiently.
Algorithm 1 Policy Gradient Algorithm for $\alpha$-Potential Distributed Game $\mathcal{G}$
1:Input: A policy class $\{\phi_{\theta}:[0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}\to A\mid\theta\in\mathbb{R}^{L}\}$, time grid $\pi_{P}$, mini-batch sample size $M\in{\mathbb{N}}$, and learning rates $(\tau_{n})_{n\geq 0}\subset(0,\infty)$.
2:Initialize: initial parameter $\theta_{0}$.
3:for $n=0,1,\ldots$ do
4:  Generate $M$ independent trajectories from (5.3) with policy $\phi_{\theta_{n}}$.
5:  Evaluate the cost $J_{M}({\theta_{n}})$ by (5.4) using the sampled trajectories.
6:  Update $\theta$: $\theta_{n+1}=\theta_{n}-\tau_{n}\nabla_{\theta}J_{M}(\theta_{n})$.
7:end for
8:Output: approximate policy $\phi_{\theta^{*}}$.
Note that at each iteration, Algorithm 1 performs a gradient descent update for all players’ policy parameters simultaneously. In comparison, the standard fictitious play algorithm (see [19]) entails a significantly higher computational cost, as it requires solving $N$ individual stochastic control problems at each iteration for each player’s best response to other players’ previous controls. Each of these sub-problems typically requires hundreds or even thousands of gradient descent updates.
The $\alpha$-potential structure of the game $\mathcal{G}$ is essential in reducing the computation of $\alpha$-NEs to the minimization of a common objective function $\Phi$. This structure is key to ensuring the convergence of the gradient-based updates in Algorithm 1. While policy gradient methods converge for various stochastic control problems (see e.g., [37, 14, 40]), it is well known that they may diverge in general multi-agent games without additional structure assumptions [31].
6. Application to Game-theoretic Motion Planning
This section illustrates our results using the crowd motion game from Section 1, which is a special case of the distributed games introduced in Section 2. These games offer an agent-based framework for modeling crowd dynamics, where each pedestrian makes rational decisions to control their motion based on individual preferences, and the resulting equilibrium behavior determines the evolution of the crowd.
Specifically, given a joint control profile ${\textbf{{u}}}=(u_{i})_{i\in[N]}\in\mathcal{H}^{2}({\mathbb{R}}^{kN})$, player $i$ considers the following objective function (cf. (1.2)):
$J_{i}({\textbf{{u}}})\coloneqq{\mathbb{E}}\left[\int_{0}^{T}\left(\ell_{i}(u_{i,t})+\frac{1}{N-1}\sum_{j=1,j\neq i}^{N}q_{ij}K(X^{u_{i}}_{i,t}-X^{u_{j}}_{j,t})\right)\,{\mathrm{d}}t+g_{i}(X^{u_{i}}_{i,T})\right],$ (6.1)
where for each $i\in[N]$, player $i$’s state process $X^{u_{i}}_{i}$ is governed by the dynamics (1.1), recalled below:
${\mathrm{d}}X_{i,t}$ $=b_{i}(t)u_{i,t}{\mathrm{d}}t+\sigma_{i}(t){\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}\gamma_{ij}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad t\in(0,T];\quad X_{i,0}=x_{i},$ (6.2)
$\ell_{i}:{\mathbb{R}}^{k}\to{\mathbb{R}}$, $K:{\mathbb{R}}^{d}\to{\mathbb{R}}$, $g_{i}:{\mathbb{R}}^{d}\to{\mathbb{R}}$ are given measurable functions, and $q_{ij}\geq 0$ is a given constant. Player $i$ aims to minimize (6.1) over the control set (see also (2.1)):
$\mathcal{A}_{i}=\{u:\Omega\times[0,T]\to A_{i}\mid u\in\mathcal{H}^{2}({\mathbb{R}}^{k}),\|u\|_{\mathcal{H}^{2}({\mathbb{R}}^{k})}\leq U\},$ (6.3)
where $U>0$ is a sufficiently large constant.
In this game, each player aims to reach their respective destination, specified by the terminal costs $(g_{i})_{i\in[N]}$, at a given terminal time, with their preferred route influenced by the spatial distribution of the population through the kernel $K$ and the interaction weights $(q_{ij})_{i,j\in[N]}$. Depending on the structure of the kernel $K$, the game can model self-organizing behavior (commonly referred to as flocking), or aversion behavior, as discussed in detail below.
Example 6.1 (Kernel choices).
When $K$ decreases as the distance between players increases, the game models congestion-averse behavior, such as pedestrians avoiding densely populated areas. One such choice is the Gaussian-type kernel
$K(z)=\exp\left(-\rho{|z|^{2}}\right),\quad\textnormal{with $\rho>0$},$ (6.4)
analogous to the exponentially decaying repulsion function used in collision-avoidance pedestrian models [44]. An alternative kernel is the following smoothed indicator function:
$K(z)\coloneqq\int_{{\mathbb{R}}^{d}}\mathbf{1}_{B_{r}}(z-v)\gamma_{\delta}(v){\mathrm{d}}v,$ (6.5)
where $\gamma_{\delta}(v)\coloneqq\frac{1}{\delta}\gamma\left(\frac{v}{\delta}\right)$ is a mollifier, with $\gamma:{\mathbb{R}}^{d}\to{\mathbb{R}}$ being a smooth function with compact support, and $\mathbf{1}_{B_{r}}$ is the indicator of the ball $B_{r}$ centered at 0 with radius $r>0$. This kernel function (6.5) has been used in the nonlocal aversion model [3], which captures the phenomenon that each pedestrian is only affected by crowding within their personal space $B_{r}$.
When $K$ increases with the distance between players, the model promotes aggregation, mimicking coordinated motion in flocks or herds, which is driven by factors such as safety, energy efficiency, or social alignment. To model such a self-organizing behavior, one may use the following quadratic kernel as in [17]:
$K(z)=\frac{1}{2}|z|^{2},$
or the Cucker–Smale-type flocking kernel used in [39].
6.1. Quantifying $\alpha$
We impose the following regularity conditions on the model coefficients.
H. 3.
For all $i,j\in[N]$, the set $A_{i}$ and the functions $b_{i},\sigma_{i}$ and $\gamma_{ij}$ satisfy (H.11). The functions $(\ell_{i})_{i\in[N]}$, $K$ and $(g_{i})_{i\in[N]}$ are twice continuously differentiable with bounded second-order derivatives.
Note that all kernel functions specified in Example 6.1 satisfy the regularity conditions in (H.3).
The following theorem specializes Theorem 3.1 to the above crowd motion game.
Theorem 6.1.
Suppose (H.3) holds. Let $B=\max_{i\in[N]}\|b_{i}\|_{L^{2}}$, and $\kappa=\|\partial^{2}_{xx}K\|_{L^{\infty}}$. The crowd motion game defined by (6.1)-(6.2) is an $\alpha_{N}$-potential game with
$\alpha_{N}\leq\frac{1}{2}TB^{2}U^{2}\frac{\kappa}{N-1}\max_{i\in[N]}\sum_{j\neq i}|q_{ji}-q_{ij}|.$ (6.6)
The upper bound of $\alpha_{N}$ in (6.6) characterizes the degree of asymmetric interactions between any two players in the dynamic game (6.1)–(6.2), expressed in terms of the time horizon, the curvature of the kernel $K$ and the interaction weights $(q_{ij})_{i,j\in[N]}$. Note that the curvature $\kappa$ can, in turn, be bounded by the parameter $\rho$ in the exponential interaction kernel (6.4), and by the parameter $r>0$ in the smoothed indicator kernel (6.5). These parameters quantify the sensitivity of each player to the distance of other players.
To derive a more explicit bound on $\alpha_{N}$, we impose additional structure on the interaction weights as follows.
  1. (a)
    Symmetric interaction. The weights $(q_{ij})_{i,j\in{[N]}}$ satisfy the pairwise symmetry condition
    $q_{ij}=q_{ji},\quad\forall i,j\in{[N]}.$ (6.7)
    This symmetry condition is satisfied when (6.1) involves mean field interactions (i.e., $q_{ij}=1$) [3, 8, 39], and more generally when the weights are derived from a symmetric graph, as in graphon mean field games (see e.g., [2]).
  2. (b)
[…]
    Suppose that $G$ has a bounded degree $\max_{i\in[N]}\deg(i)={d}_{G}$ for some ${d}_{G}\geq 2$, i.e., each player is connected to at most ${d}_{G}$ players. Additionally, we assume that the asymmetry in interactions diminishes as the graph distance between players increases. In particular, we consider the case where the degree of asymmetry exhibits an exponential decay:
    $|q_{ij}-q_{ji}|\leq w_{i,j}{\rho}^{c(i,j)},\quad\forall i,j\in[N],i\neq j$ (6.8)
    where $(w_{i,j})_{i,j\in{[N]}}$ are distinct positive constants that are uniformly bounded in $N$, $\rho\in(0,1)$ is a given constant, and $c(i,j)$ is the (shortest-path) distance between vertices $i$ and $j$. We also consider the case where the degree of asymmetry exhibits a polynomial decay:
    $|q_{ij}-q_{ji}|\leq w_{i,j}\frac{1}{c(i,j)^{\beta}},\quad\forall i,j\in{[N]},i\not=j,$ (6.9)
    where $\beta>0$ is a given constant, and $(w_{i,j})_{i,j\in{[N]}}$ are distinct positive constants that are uniformly bounded in $N$.
The following corollary refines the upper bound on $\alpha_{N}$ in Theorem 6.1 for both cases a and b, providing an explicit dependence on the number of players $N$, as well as on the parameters $\rho,d_{G}$ and $\beta$, which capture the strength and asymmetry of player interactions.
Corollary 6.1.
Suppose (H.3) holds. The crowd motion game defined by (6.1)–(6.2) is an $\alpha_{N}$-potential game with
$\alpha_{N}\leq\frac{1}{2}\kappa TB^{2}U^{2}\zeta_{N},$
where $\kappa$ and $B$ are defined as in Theorem 6.1, and $\zeta_{N}$ is determined by the structure of the interaction weights $(q_{ij})_{i,j\in[N]}$ as follows:
  1. (a)
    If $(q_{ij})_{i,j\in[N]}$ satisfies the symmetry condition (6.7), then $\zeta_{N}=0$, i.e., the game is a potential game.
  2. (b)
    If $(q_{ij})_{i,j\in[N]}$ satisfies the exponential decay condition (6.8), then as $N\to\infty$,
    $\zeta_{N}=\begin{cases}\mathcal{O}\left(N^{\frac{\ln\rho}{\ln d_{G}}}\right),&\textnormal{if }\rho\in(1/d_{G},1),\\ \mathcal{O}\left(\dfrac{\ln N}{N}\right),&\textnormal{if }\rho=1/d_{G},\\[6.0pt] \mathcal{O}\left({N}^{-1}\right),&\textnormal{if }\rho\in(0,1/d_{G}).\end{cases}$
  3. (c)
    If $(q_{ij})_{i,j\in[N]}$ satisfies the power-law decay condition (6.9), then as $N\to\infty$,
    $\zeta_{N}=\mathcal{O}\left(\dfrac{\ln\ln N}{\left(\ln N\right)^{\beta}}\right).$
Proof.
Let $\zeta_{N}=\frac{1}{N-1}\max_{i\in[N]}\sum_{j\neq i}|q_{ji}-q_{ij}|$. It is clear that $\zeta_{N}=0$ under Condition (3.14), which proves Item a. To prove Items b and c, we assume without loss of generality that for all $i,j\in[N]$ with $i\not=j$, $c(i,j)<\infty$, since otherwise $|q_{ij}-q_{ji}|=0$ under Condition (6.8) or Condition (6.9).
We first introduce the following rebalancing technique for the underlying graph $G$: Fix node $i\in[N]$. Let $T_{1}\subset G$ be the tree with node $i$ as its root. $T_{1}$ contains the shortest path for each $j\neq i$ to the root $i$, and denote $c_{1}$ by the shortest-path distance in $T_{1}$, which satisfies
$c_{1}(i,j)=c(i,j),\quad\forall j\neq i.$
We will rebalance the tree $T_{1}$ as follows to obtain a $d_{G}$-ary tree $T_{2}$, in which every node except those at the deepest level has exactly $d_{G}$ children: starting from a node $j$ that is farthest from the root, we traverse the tree (e.g., depth-first search or breadth-first search) to move $j$ to a higher level that is available, reducing its distance to the root $i$. We repeat this process until no further adjustment can be made. We denote $L+1$ as the number of levels in $T_{2}.$ Specifically, $L$ is the smallest integer that $1+d_{G}+d_{G}^{2}+\cdots+d_{G}^{L}\geq N$. So as $N\to\infty$, $N=\mathcal{O}(d_{G}^{L})$ and $L=\mathcal{O}(\frac{\ln N}{\ln d_{G}}).$ Let $c_{2}$ denote the distance in $T_{2}$. Since the reblancing process shortens the distance between the nodes,
$c_{2}(i,j)\leq c_{1}(i,j)=c(i,j),\quad j\not=i.$
For Item b, there exists a constant $C\geq 0$, which depends only on $(w_{ij})_{i,j\in[N]}$ and $d_{G}$, such that
$\zeta_{N}\leq\frac{C}{N}\max_{i\in[N]}\sum_{j\neq i}\rho^{c(i,j)}\leq\frac{C}{N}\sum_{\ell=1}^{L}\rho^{\ell}d_{G}^{\ell},$ (6.10)
where the first inequality follows from Condition (6.8), and the last inequality uses $\rho\in(0,1)$ and the rebalancing technique, which is an upper bound of the summation of weights in $T_{2}$. It remains to compute the right-hand side of (6.10). If $\rho d_{G}=1$,
$\zeta_{N}\leq\frac{CL}{N}=\mathcal{O}\left(\frac{\ln N}{N\ln d_{G}}\right).$ (6.11)
If $\rho d_{G}\neq 1$,
$\zeta_{N}\leq\begin{cases}C\frac{1}{d_{G}^{L}}\rho d_{G}\frac{(\rho d_{G})^{L}-1}{\rho d_{G}-1}\leq C\rho^{L}=\mathcal{O}\left(N^{\frac{\ln\rho}{\ln d_{G}}}\right),&\textnormal{if }\rho d_{G}>1,\\ \frac{C}{N}\frac{\rho d_{G}}{1-\rho d_{G}}=\mathcal{O}\left(N^{-1}\right),&\textnormal{if }\rho d_{G}<1.\end{cases}$ (6.12)
Combining (6.11) and (6.12) finishes the proof for Item b.
For Item c, fix $i\in[N]$, let $n_{\ell}$ denote the number of nodes at distance $\ell$ from the root in $T_{1}$. Then under Condition (6.9),
$\zeta_{N}\leq\frac{2}{N}\sum_{j\not=i}|q_{ij}-q_{ji}|\leq\frac{2}{N}\left(\max_{i,j\in[N]}|w_{ij}|\right)\sum_{\ell=1}^{N}\frac{n_{\ell}}{\ell^{\beta}}\leq 2\left(\max_{i,j\in[N]}|w_{ij}|\right)\frac{1}{N}\sum_{\ell=1}^{L}\frac{d_{G}^{\ell}}{\ell^{\beta}},$ (6.13)
where the last inequality provides an upper bound of $\sum_{\ell=1}^{N}\frac{n_{\ell}}{\ell^{\beta}}$ using the rebalanced tree $T_{2}$. Observe that the function $h(x)\coloneqq d_{G}^{x}/x^{\beta}$ has the derivative $h^{\prime}(x)=\frac{d_{G}^{x}(x\ln d_{G}-\beta)}{x^{\beta+1}}$, and is increasing on $(\beta/\ln d_{G},\infty)$. Hence for all $M\in\{1,\ldots,L\}$ with $M\geq\beta/\ln d_{G}$,
$\sum_{\ell=1}^{L}\frac{(d_{G})^{\ell}}{\ell^{\beta}}\leq\sum_{\ell=1}^{L-M}\frac{(d_{G})^{\ell}}{\ell^{\beta}}+M\frac{(d_{G})^{L}}{L^{\beta}}$ (6.14)
Since $x\to 1/x^{\beta}$ is deceasing on (0,)(0$\infty$)( 0 , ∞ ), the first term on the right-hand side of (6.14) can be upper bounded by
$\sum_{\ell=1}^{L-M}\frac{1}{\ell^{\beta}}\leq 1+\int_{1}^{L-M}\frac{1}{x^{\beta}}{\mathrm{d}}x=\begin{cases}1+\frac{1}{\beta-1}\left(1-{(L-M)}^{(1-\beta)}\right),&\text{ if }\beta>1,\\ 1+\ln(L-M),&\text{ if }\beta=1,\\ 1+\frac{1}{1-\beta}\left({(L-M)}^{(1-\beta)}-1\right),&\text{ if }0<\beta<1.\end{cases}$
Thus for $\beta>1$, taking $M^{*}=\beta\left\lfloor\frac{\ln L}{\ln{d_{G}}}\right\rfloor$, which implies that $(d_{G})^{M^{*}}=\mathcal{O}(L^{\beta})$ as $L\to\infty$. By (6.14),
$\sum_{\ell=1}^{L}\frac{(d_{G})^{\ell}}{\ell^{\beta}}\leq C\left((d_{G})^{L-M^{*}}+\ln L\frac{(d_{G})^{L}}{L^{\beta}}\right)\leq C\ln L\frac{(d_{G})^{L}}{L^{\beta}},$ (6.15)
which along with (6.13) shows that as $N\to\infty$,
$\zeta_{N}\leq C\frac{1}{N}\ln L\frac{(d_{G})^{L}}{L^{\beta}}\leq C\ln\ln N\left(\frac{1}{\ln N}\right)^{\beta}.$
For $\beta=1$, taking $M^{*}=\left\lfloor\frac{\ln L}{\ln{d_{G}}}\right\rfloor$, which implies that $(d_{G})^{M^{*}}=\mathcal{O}(L)$ as $L\to\infty$. By (6.14),
$\sum_{\ell=1}^{L}\frac{(d_{G})^{\ell}}{\ell^{\beta}}\leq C\left((d_{G})^{L-M^{*}}\ln L+\ln L\frac{(d_{G})^{L}}{L}\right)\leq C\ln L\frac{(d_{G})^{L}}{L},$ (6.16)
which along with (6.13) implies $\zeta_{N}=\mathcal{O}\left(\ln\ln N\left(\frac{1}{\ln N}\right)\right)$ as $N\to\infty$. Similarly, for $\beta\in(0,1)$, taking $M^{*}=\left\lfloor\frac{\ln L}{\ln{d_{G}}}\right\rfloor$ and using (6.14) yield
$\sum_{\ell=1}^{L}\frac{(d_{G})^{\ell}}{\ell^{\beta}}\leq C\left((d_{G})^{L-M^{*}}L^{1-\beta}+\ln L\frac{(d_{G})^{L}}{L^{\beta}}\right)\leq C\ln L\frac{(d_{G})^{L}}{L^{\beta}},$ (6.17)
which along with (6.13) implies $\zeta_{N}=\mathcal{O}\left(\ln\ln N\left(\frac{1}{\ln N}\right)^{\beta}\right)$ as $N\to\infty$. This completes the proof. ∎
6.2. Numerical results for NEs
We apply Algorithm 1 to compute the NEs in the crowd motion game (6.1)–(6.2). For ease of exposition, we consider a four-player game (i.e., $N=4$), where each player has two-dimensional state and control processes (i.e., $d=k=2$ and $A_{i}={\mathbb{R}}^{2}$). Player $i$’s state dynamics is given by
${\mathrm{d}}X_{i,t}^{u_{i}}$ $=u_{i,t}{\mathrm{d}}t+\sigma_{i}{\mathrm{d}}W^{i}_{t}+\gamma_{i}{\mathrm{d}}\widetilde{\eta}_{i,t}+\gamma_{0}{\mathrm{d}}\widetilde{\eta}_{0,t},\quad X_{i,0}=x_{i,0},$ (6.18)
where $\sigma_{i},\gamma_{i},\gamma_{0}\geq 0$ are given constants, $W^{i}$ an $\tilde{\eta}_{i}$ are two-dimensional Brownian motion and compensated Poisson processes, respectively, representing the idiosyncratic noise for player $i$, and $\tilde{\eta}_{0}$ is an independent two-dimensional compensated Poisson process modeling the common noise shared by all players. The process $\tilde{\eta}_{i}$ has a constant intensity $\lambda_{i}$, with $\lambda_{0}=0.25$, $\lambda_{1}=0.3$, and $\lambda_{i}=0.2$ for all $i\geq 2$. Player $i$ considers minimizing the objective (6.1) with the terminal time $T=1$, and terminal cost
$g_{i}(x)=c_{i}|x-z_{i}|^{2},$ (6.19)
where $c_{i}>0$ is a given constant, and $z_{i}\in{\mathbb{R}}^{d}$ is the target that player $i$ aims to reach at time $T$. The running cost $\ell_{i}$, the kernel $K$ and the interaction weights $(q_{ij})_{i,j=1}^{N}$ will be specified below. Algorithm 1 is implemented using neural network-based policies, with the detailed architecture and training procedures described in Appendix A.
6.2.1. Aversion Games with Idiosyncratic Noises.
We first consider a crowd-aversion game in which all players are subject only to idiosyncratic noise. Specifically, we set $\sigma_{i}=0.1(i-1)/N$, $\gamma_{i}=0.1$, and $\gamma_{0}=0$ in (6.18). All players start from the same initial location $x_{i,0}=(0,0)$, and aim to reach a common terminal location $z_{i}=(0.5,0.5)$. The terminal cost function $g_{i}$ is given by (6.19) with $c_{i}=1$, and the running cost $\ell_{i}$ on control is $\ell_{i}(a)=\frac{0.1}{2}|a|^{2}$. To model crowd-aversion effects, we adopt the Gaussian kernel $K(z)=100\exp(-100|z|^{2})$, and assume uniform interaction weights $q_{ij}=1$ in (6.1), representing symmetric aversion among all players. The resulting crowd motion game is a potential game as shown in Corollary 6.1.
Figure 1 illustrates the equilibrium trajectories of the players, where positions at times $t=0.25,0.5,0.75$ are marked by symbols 1, 2, and 3, respectively. The left panel shows the mean positions computed over 500500500 sample trajectories, while the right panel presents a representative single-sample trajectory.
All players begin at the same initial location (indicated by a red circle at position (0,0)(0,0)( 0 , 0 )) and aim to reach a common target (marked by a red cross at (0.5,0.5)(0.5,0.5)( 0.5 , 0.5 )). Early in the game, players disperse in different directions to reduce crowding, a behavior induced by the pairwise aversion term in the cost function. Notably, Player 4 takes a wide detour to avoid other players before converging near the destination. The group exhibits loose coordination: although all players share the same goal, their individual trajectories reflect mutual avoidance dynamics.
Figure 1. Equilibrium trajectories in the aversion game with a Gaussian kernel and uniform interaction weights. Left: mean positions over 500 simulations. Right: one representative trajectory. The solid circle denotes the shared initial position; the cross marks the common target. Markers “1”, “2”, and “3” indicate positions at times 0.250.250.25, 0.50.50.5, and 0.750.750.75, respectively.
6.2.2. Flocking Games with Idiosyncratic Noises.
The second example considers a flocking game where all players start from the same initial location $x_{i,0}=(0,0)$, and aim for distinct individual target: (0.25,0)(0.25,0)( 0.25 , 0 ), (0,0.5)(0,0.5)( 0 , 0.5 ), (0.5,0)(-0.5,0)( - 0.5 , 0 ), and (0,1)(0,-1)( 0 , - 1 ). Each player is influenced only by idiosyncratic noise, with parameters set as $\sigma_{i}=0.1(i-1)/N$, $\gamma_{i}=0.1$, and $\gamma_{0}=0$ in (6.18). The flocking behavior is modeled using the quadratic kernel $K(z)=\frac{1}{2}|z|^{2}$. Each player $i$ incurs a running cost on control given by $c_{i}(a)=\frac{0.1}{2}|a|^{2}$, and a terminal cost defined by (6.19), with $c_{i}=40$.
We consider two different settings for the interaction weights $(q_{ij})_{i,j=1}^{4}$ in (6.1). In the first setting, uniform interaction is assumed, with $q_{ij}=1$ for all $i\neq j$, so that each player is equally influenced by every other player. In the second setting, a two-group structure is imposed: players 2 and 3 form one group, and players 1 and 4 form another. In this case, $q_{ij}=1$ if players $i$ and $j$ belong to the same group, and $q_{ij}=0$ otherwise. This models selective flocking behavior, where players tend to coordinate only with those in their own group.
Figure 2 shows the equilibrium trajectories under uniform interaction weights. In this case, the group first aggregates toward a common intermediate point and, after time $t=0.5$, the players begin to diverge toward their individual destinations. In contrast, Figure 3 presents the equilibrium trajectories under the two-group interaction structure. Here, each subgroup converges toward a distinct intermediate point, illustrating that the interaction structure encoded in $(q_{ij})_{i,j=1}^{N}$ has a significant impact on both the alignment dynamics and the overall configuration of the players.
Figure 2. Equilibrium trajectories in the flocking game with a quadratic kernel and uniform interaction weights. Left: mean positions over 500 simulations. Right: one representative trajectory. The solid circle denotes the shared initial position; the crosses mark the individual targets. Markers “1”, “2”, and “3” indicate positions at times 0.25, 0.5, and 0.75, respectively.
Figure 3. Equilibrium trajectories in the flocking game with a quadratic kernel and group-based interaction weights. Players 1 and 4 belong to one group, and players 2 and 3 form the other. The interaction weights $q_{ij}=1$ if players $i$ and $j$ are in the same group, and $q_{ij}=0$ otherwise. Left: mean positions over 500 simulations. Right: one representative trajectory. The solid circle denotes the shared initial position; the crosses mark the individual targets. Markers “1”, “2”, and “3” indicate positions at times 0.25, 0.5, and 0.75, respectively.
6.2.3. Flocking Games with Common Noises.
To demonstrate the flexibility of our framework, we consider a flocking game driven solely by common jumps. Specifically, we set $\sigma_{i}=\gamma_{i}=0$, and $\gamma_{0}=0.1$, so that only common noise influences the dynamics. All other model parameters are identical to those in the previous flocking game with uniform interaction weights.
Figure 4 presents two sample trajectories of the resulting equilibrium dynamics. The common jumps introduce abrupt, synchronized shifts in the players’ positions, followed by realignment as they continue moving toward their respective targets. While the jump events cause irregularities in the intermediate paths, the overall flocking behavior remains consistent with the patterns observed in the present setting with pure idiosyncratic noises.
Figure 4. Equilibrium trajectories in the flocking game with a quadratic kernel, uniform interaction weights, and pure common jumps. The solid circle denotes the shared initial position; the crosses mark the individual targets. Markers “1”, “2”, and “3” indicate positions at times 0.25, 0.5, and 0.75, respectively.
Appendix A Implementation of Algorithm 1 for Crowd-Motion Games
To implement Algorithm 1, we uniformly discretize the time interval [0,1][0,1][ 0 , 1 ] into $L=50$ steps. The batch size $M$, representing the number of simulated trajectories per parameter update, is set to 500.
Before stating the configuration details of policy parameterisation, we remark that the algorithm’s hyperparameters have not been optimally tuned and hence the following choices may not represent the optimal combination.
We employ a residual feedforward neural network architecture following [28], consisting of an input layer, a sequence of residual blocks, and an output layer. Each residual block has the form $x\mapsto\varrho(L_{1}(\varrho(L_{2}(x))))+x$ where $L_{1}$ and $L_{2}$ are fully connected layers with matching input and output dimensions, and $\varrho$ denotes the activation function, chosen here to be the standard ReLU.
Our neural network policies comprise four residual blocks, each with width $d+10$, where $d=4\times 4+1=17$ is the dimensions of the joint state and sensitivity processes, and also the time variable. This neural architecture requires no prior knowledge of the solution’s structure. Parameters are optimized using the Adam optimizer, with an initial learning rate of $10^{-3}$. A ReduceLROnPlateau scheduler from PyTorch is employed to automatically reduce the learning rate when the validation loss stagnates. All experiments are run using the fixed random seed 2025.
All experiments are conducted on a MacBook Pro with 16GB of memory and a Apple M1 Pro chip.